OpenStudy (anonymous):

Simplify the following expression and write the answer with positive exponents. 12m^4 - 24m^3 + 6m^2 + 9m-9 / 6m^2

OpenStudy (mertsj):

\[\frac{12m^4}{6m^2}-\frac{24m^3}{6m^2}+\frac{6m^2}{6m^2}+\frac{9m ^{-9}}{6m^2}=\] \[2m^2-4m+1+\frac{3}{2m ^{11}}\]